# Initialization is Critical to Whether Transformers Fit Composite Functions by Inference or Memorizing

### v2

@@ Appendix A Experimental Setups
For Fig. 2(c), we employ the architecture of GPT-2-SMALL in our experiments. Specifically, to ensure consistency between the model structure and the settings described in the paper, we adopt the post-norm configuration. We have also conducted experiments using the pre-norm setting and obtained consistent conclusions. Apart from the differences in the model architecture, we retain the same data, training strategy, parameter initialization, and other settings as mentioned in the previous sections.
For Fig. 6, most of the experimental settings are the same. The difference lies in the dataset size. Specifically, we train the transformer models on a dataset of 100,000 samples. To increase the inferential complexity, we systematically modify the designated target mappings of the following anchor pair groups: {(1, 2), (2, 1)}, {(1, 3), (3, 1)}, {(1, 4), (4, 1)}, {(2, 3), (3, 2)}, {(2, 4), (4, 2)}, and {(3, 4), (4, 3)}.
@@ E.1 Detailed Results for Condensation with Different Learning Rates and Weight Decay Coefficients
Appendix F The fitting speed of different types of anchor pairs
@@
In this section, we demonstrate the fitting speed of different anchor pairs when the inferential complexity is one. In this experiment, we simultaneously set the anchor pairs (1, 2) and (2, 1) as the same non-inferential mappings. As shown in Fig. 16, we present the number of epochs required for 16 types of anchor pairs to achieve 60% accuracy on their corresponding training data. We highlight the four types of anchor pairs (1, 2), (2, 1), (1, 1), and (2, 2) in red (since we set (1, 2) and (2, 1) as non-inferential mappings, the above four types of anchor pairs will be significantly affected). It can be observed that the highlighted anchor pair types exhibit notably slower training speeds, further verifying that the presence of non-inferential mappings reduces the training speed of models with small initialization.

### v3

@@ Appendix A Experimental Setups
For Fig. 2(c), Fig. 6, Fig. 7, Fig. 17 we employ the architecture of GPT-2 in our experiments. For Fig. 2(c), to ensure consistency between the model structure and the settings described in the paper, we adopt the post-norm configuration. We have also conducted experiments using the pre-norm setting and obtained consistent conclusions. Apart from the differences in the model architecture, we retain the same data, training strategy, parameter initialization, and other settings as mentioned in the previous sections.
@@
For Fig. 18, Fig. 19 most of the experimental settings are the same. The difference lies in the dataset size. Specifically, we train the transformer models on a dataset of 100,000 samples. To increase the inferential complexity, we systematically modify the designated target mappings of the following anchor pair groups: {(1, 2), (2, 1)}, {(1, 3), (3, 1)}, {(1, 4), (4, 1)}, {(2, 3), (3, 2)}, {(2, 4), (4, 2)}, and {(3, 4), (4, 3)}.
@@ E.1 Detailed Results for Condensation with Different Learning Rates and Weight Decay Coefficients
Appendix F Detailed Introduction of Datasets and Additional Results
Compositional tasks: SCAN and COGS. SCAN and COGS are two classic compositional tasks, both of which are also synthetic but with more natural language variance, thus being ideal testbeds. For the SCAN dataset, we selected the “Generalizing composition across primitive commands" task, where the “turn left" command only appears in single-command mappings and is trained alongside other composite commands. We assess the model’s generalization ability on composite commands that include the “turn left" command. For the COGS dataset, we evaluate the model’s in-distribution and out-of-distribution generalization performance after training on the same training set. In-distribution generalization is tested with data constructed from different primitives in the same combinatorial patterns. Out-of-distribution generalization is tested with data that follows different combinatorial rules (the original paper outlines 21 methods for generating out-of-distribution data, from which we generate test data with equal probability).
Reasoning tasks: PrOntoQA. PrOntoQA is a synthetic multi-step reasoning dataset where each data point assigns hierarchical relationships among objects and requires the model to determine whether a multi-step reasoning chain is correct. Although we use next-token prediction for training, during testing, we only evaluate the model’s accuracy in judging hierarchical relationships (thus, the model’s random guessing accuracy is 50%).
Realistic tasks: Addition task and SlimPajama dataset. Unlike traditional addition tasks, we use a case-based reasoning intervention experiment (Hu et al., 2024) to study the generalization of rule learning in the addition task. Specifically, we consider the setup: $a+b=c$, where $a,b\in[0,999]$. We use $a,b\in[400,600]$ as the test set and the remaining data as the training set. This construction prevents the model from simply mimicking training data similar to the test set. We trained a simple 2-layer 1-head model and a GPT-2 model. As shown in Fig. 17(a), regardless of model size and learning mode, small initialization scales (or large weight decay coefficients) generally lead to good rule generalization, while large initialization scales (or small weight decay coefficients) fail to generalize perfectly.
Figure 17: Performance comparison of models with different initialization scales and weight decay coefficients on realistic tasks: Addition task and SlimPajama dataset. (a) In the addition task, we use a case-based reasoning intervention experiment with a test set of $a,b\in[400,600]$ and the remaining data as the training set. Models with small initialization scales (or large weight decay) generally show better rule generalization. (b) For the SlimPajama dataset, GPT-2 Medium models trained on GitHub and GitHub+Wikipedia data with small initialization scales consistently achieved lower perplexity. The parameters are initialized following a zero-mean normal distribution with a standard deviation of $d_{in}^{-\gamma}$.
For the SlimPajama dataset, we used two data compositions: the GitHub section and the GitHub+Wikipedia section. We trained GPT-2 Medium models with different initializations on both datasets for 40B tokens. As shown in Fig. 17(b), for both data compositions, smaller initialization scales consistently achieved lower perplexity. Notably, by observing the training trajectories, we found that for the GitHub data, the model with small initialization achieved lower perplexity than the model with large initialization early in training (around 2B tokens). For the GitHub+Wikipedia data, the model with small initialization achieved lower perplexity later in training (around 30B tokens). This further validates the preference of small initialization for reasoning tasks. We will present the detailed training trajectories in the revised manuscript.
Appendix G Prediction of Mechanisms underlying Different Solutions
To measure the training steps required to fit data, we use the number of epochs required for the model’s training loss to reach $5\times 10^{-2}$. We conduct 9 random trials for each setting. As shown in Fig. 18, models that tend to learn data with inferential solutions indeed use more training steps for data with larger inferential complexity (Fig. 18(a)), while models that tend to learn data with symmetric solutions use roughly the same training steps for data with different inferential complexity (Fig. 18(b)). This prediction convincingly backs up our analysis of why different initialization scales can lead to different solutions.
@@ Appendix G Prediction of Mechanisms underlying Different Solutions
We further demonstrate the fitting speed of different anchor pairs when the inferential complexity is one. In this experiment, we simultaneously set the anchor pairs (1, 2) and (2, 1) as the same non-inferential mappings. As shown in Fig. 19, we present the number of epochs required for 16 types of anchor pairs to achieve 60% accuracy on their corresponding training data. We highlight the four types of anchor pairs (1, 2), (2, 1), (1, 1), and (2, 2) in red (since we set (1, 2) and (2, 1) as non-inferential mappings, the above four types of anchor pairs will be significantly affected). It can be observed that the highlighted anchor pair types exhibit notably slower training speeds, further verifying that the presence of non-inferential mappings reduces the training speed of models with small initialization.